# On characterization and existence of a constrained correlated equilibria in Markov games

### v1

@@ A.3 Main proofs of Section 3
=$\overset{(i)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}^{i}}\sum_{\hat{a}_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\bar{r}_{t}^{i,\pi}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})$
=$\overset{(ii)}{=}$ $\sum_{t=1}^{H}|\mathcal{A}^{i}|^{t}\sum_{s_{t},a_{t}^{i}}\sum_{\hat{a}_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\left(\sum_{a_{t}^{-i}}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t})\right)$
[…]
=$\overset{(iii)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t}}\sum_{(\hat{a}_{t}^{i},a_{t}^{-i})}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\left(d_{t}^{\bar{\phi}^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\right)$
@@ B.1 Verification of Example 1
    A constrained correlated equilibrium does not exist: Given any feasible policy $\pi$, the best $i$-feasible modification $\phi^{2}\in\Phi^{2}_{NM}$ for P2, $i=2$, is to choose $\phi^{2}(\tilde{a}^{2}|a^{2})=\mathds{1}_{\tilde{a}^{2}=2}$ for all $a^{2}\in\{1,2\}$ (i.e., always modify to play $a^{2}=2$), as this maximizes $V^{r^{2}}(\pi)$ while satisfying the constraint.
    $P1$ $P2$ $1$ $2$
    $1$    0 y+x
    $2$ 0 1-x-y
    However, this new policy $\phi^{2}\circ\pi$ is not $i$-feasible where $i=1$ since P1’s constraint is violated.
@@ B.2 Verification of Example 2
    We first verify part (a) in the second condition of Assumption 4.2:
    $V^{g^{j}}(\phi^{i}(3)\circ\pi)=0<\frac{1}{4},\quad j=1,2,$
    $V^{g^{j}}(\phi^{i}(2)\circ\pi)=0<\frac{1}{4},\quad j=3,4.$
    Next, we verify part (b) in the second condition of Assumption 4.2, namely if we choose the vector $\alpha=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$, we have:
@@
    For P2 the second condition of Assumption 4.2 is satisfied by analogous arguments. Thus, Assumption 4.2 is satisfied.
@@ C.1.1 Sensitivity analysis of $LP^{i}(\pi)$
$\max_{\alpha\in\mathbb{R}^{K^{i}}}$ $\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)$
@@
We furthermore consider the following assumptions:
@@
There exists a strictly feasible solution $\alpha\in\Delta(K^{i})$ such that
$\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)>c^{j},\,\forall j\in[J].$
Assumption C.4.
Assume the following hold:
@@
    For all $j\in[J]$ it holds that $\not\exists\beta\in\mathbb{R}$ such that $\{V^{g^{j}}(\phi^{i}(k)\circ\pi)\}_{k=1}^{K^{i}}=\beta\mathbf{1}_{K^{i}}$.
@@
Let Assumption C.2 hold. Furthermore, let either Assumption C.2 or Assumption C.4 hold. Assume that $\pi_{n}\to\pi$ pointwise as $n\to\infty$, then the following is satisfied: [^4] Pointwise convergence of $\pi_{n}$ to $\pi$ ensures that for any $i\in\mathcal{N}$, $j\in[J]$, and $k\in[K^{i}]$: $V^{r^{i}}(\phi^{i}(k)\circ\pi_{n})\to V^{r^{i}}(\phi^{i}(k)\circ\pi)\quad\text{and}\quad V^{g^{j}}(\phi^{i}(k)\circ\pi_{n})\to V^{g^{j}}(\phi^{i}(k)\circ\pi)$ pointwise as $n\to\infty$.
@@
    Consider $rank(\psi)$ and $rank(\psi_{n})$: Since $\psi_{n}=\psi=\sum_{k=1}^{K^{i}}\alpha_{k}-1$, therefore
@@
    Case I (Assumption C.3 holds) Under Assumption C.3 there exists a strictly feasible solution $\alpha$ which implies that $\xi_{j}(\alpha)<0$ for all $j\in[J]$, and $\psi(\alpha)=0$.
    Next, pick any solution $\alpha^{\prime}$ such that $\alpha^{\prime}_{k}>0$ for all $k\in[K^{i}]$. Then, let $\alpha^{{}^{\prime\prime}}=(1-\varepsilon)\alpha+\varepsilon\alpha^{\prime}$. For $\alpha^{{}^{\prime\prime}}$ it holds that $\psi(\alpha^{{}^{\prime\prime}})=0$ since $\psi$ is affine, and $\alpha^{{}^{\prime\prime}}_{k}>0$ for all $k\in[K^{i}]$. Furthermore, by choosing $\varepsilon$ sufficiently small, it holds that $\sum_{k=1}^{K^{i}}\alpha^{{}^{\prime\prime}}_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)>c^{j}$ and therefore $\xi_{k}(\alpha)<0$ for all $k\in\{J+1,\dots,J+K^{i}\}$.
    This implies that $\xi_{m}(\alpha^{{}^{\prime\prime}})<0$ for all $m\in[J+K^{i}]$ and therefore $I$ is empty. We conclude that:
    $rank(\xi_{n,I},\psi_{n})\overset{(i)}{=}rank(\psi_{n})\overset{(ii)}{\leq}rank(\psi)\overset{(iii)}{=}rank(\phi_{I},\psi),\,\forall n,$
    where in $(i)$ and $(iii)$ we use that $I=\emptyset$, and in $(ii)$ we use that $\psi$ has full rank. Thus, Inequality (28) is satisfied.
    Case II (Assumptions C.4 hold) By second condition of Assumption C.4, there exists a $\alpha\in\Delta(K^{i})$ such that $\alpha\in H(\xi,\psi)$, and satisfies $\alpha_{k}>0$ for all $k\in[K^{i}]$. We conclude that $\xi_{m}(\alpha)<0$ for all $m=J+1,\dots,J+K^{i}$ and thus $\{J+1,\dots,J+K^{i}\}\notin I$.
    Without loss of generality, we assume that $rank(\{\xi_{j}\}_{j\in[J]})=J$ since otherwise the redundant constraints can be removed. Note that removing redundant constraints from $LP^{i}(\pi)$ does not affect $LP^{i}_{n}$ since $I=\{m\in[J+K^{i}]\ \lvert\ \xi_{m}(\alpha)=0,\,\forall\alpha\in H(\xi,\psi)\}$ depends only on the limiting $LP^{i}(\pi)$. Furthermore, without loss of generality, we assume that $J=1$ and drop the index $j$ from our notation. For all feasible solutions, we assume that the constraint holds with equality such that $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g}(\phi^{i}(k)\circ\pi)=c$ for all $\alpha\in H(\xi,\psi)$. This implies that $\xi(\alpha)=0$ for all $\alpha\in H(\xi,\psi)$. Suppose that $(\xi_{I},\psi)$ has full rank, i.e., $rank(\xi_{I},\psi)=2$, then the following holds:
    $rank(\xi_{n,I},\psi_{n})\leq rank(\xi_{I},\psi),\,\forall n$
    since $rank(\xi_{n,I},\psi_{n})\leq 2$. Thus, Inequality (28) is satisfied.
    Next, we show that $rank(\xi_{I},\psi)=2$. Any linear mapping such as $(\xi_{I},\psi):\mathbb{R}^{K^{i}}\rightarrow\mathbb{R}^{2}$ can be represented as $\alpha^{T}M-(1,c)^{T}$ where $M\in\mathbb{R}^{2\times K^{i}}$ and is defined as follows:
    $M:=\begin{bmatrix}1,1,\dots,1\\ V^{g}(\phi^{i}(1)\circ\pi),V^{g}(\phi^{i}(2)\circ\pi),\dots,V^{g}(\phi^{i}(K^{i})\circ\pi)\end{bmatrix}$
    Therefore, by the first condition of Assumption C.4, we have $rank(M)=2$.
    Both Case I and Case II satisfy Inequality (28), which concludes the proof of the first point.
@@
    This follows from the first condition of Assumption 4.2 together with Lemma A.8.
    Assumption C.4:
    By part (a) of the second condition of Assumption 4.2 together with Lemma A.8, we know that for any $j\in[J]$ and any policy $\pi$ lying on the boundary of $\mathcal{C}_{\pi}$, there exists a Markovian deterministic modification $\phi^{i}(k)\in\Phi^{i}_{\mathrm{M,det}}$ such that $V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j}$. Meanwhile, under the identity modification $\phi^{i}(\mathrm{id})\in\Phi^{i}_{\mathrm{M,det}}$, we have $V^{g^{j}}(\phi^{i}(\mathrm{id})\circ\pi)=c^{j}$. Therefore, the first condition of Assumption C.4 follows directly from part (a) of the second condition of Assumption 4.2.
    The second condition of Assumption C.4 is satisfied by part (b) of the second condition in Assumption 4.2.
@@
=$\overset{(i)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}}\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\phi^{i}(k)\circ\pi}(s_{t},a_{t})g_{t}^{j}(s_{t},a_{t})$

### v2

@@ A.3 Main proofs of Section 3
$\overset{(i)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}^{i}}\sum_{\hat{a}_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\bar{r}_{t}^{i,\pi}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})$
$\overset{(ii)}{=}$ $\sum_{t=1}^{H}|\mathcal{A}^{i}|^{t}\sum_{s_{t},a_{t}^{i}}\sum_{\hat{a}_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\left(\sum_{a_{t}^{-i}}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t})\right)$
${=}$ $\sum_{t=1}^{H}\sum_{s_{t}}\sum_{(\hat{a}_{t}^{i},a_{t}^{-i})}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\left(|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})\right)$
$\overset{(iii)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t}}\sum_{(\hat{a}_{t}^{i},a_{t}^{-i})}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\left(d_{t}^{\bar{\phi}^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\right)$
@@ B.1 Verification of Example 1
    A constrained correlated equilibrium does not exist: By Definition 3.1, a constrained correlated equilibrium does not exist if for any feasible policy, there exists a player $i$ with an $i$-feasible modification that strictly increases player $i$’s utility. Given any feasible policy $\pi$, player 2’s best modification is
    $\phi^{2}(\hat{a}^{2}|a^{2})=\mathbf{1}_{\{\hat{a}^{2}=2\}},$
    namely, always modifying to action $a^{2}=2$. Such modification results in the policy of the form:
    $P1$ $P2$ 11 22
    11   0 y+x
    22 0 1-x-y
    This modification is always $2$-feasible and strictly increases player 2’s reward.
    Therefore, for every feasible policy, player 2 admits a $2$-feasible modification that improves its reward. Hence, no constrained correlated equilibrium exists. However, this new policy $\phi^{2}\circ\pi$ is not $i$-feasible where $i=1$ since P1’s constraint is violated.
@@ B.2 Verification of Example 2
    Now, we choose the vector $\alpha=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$, we have:
@@
    For P2, Assumption 4.2 is satisfied by analogous arguments. Thus, Assumption 4.2 is satisfied.
@@ C.1.1 Sensitivity analysis of $LP^{i}(\pi)$
$\max_{\alpha\in\mathbb{R}^{K^{i}}}\$$\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)$
@@
We furthermore consider the following assumption:
@@
Assume the following holds:
@@
    For all $j\in[J]$, if there exists a strictly infeasible $\alpha\in\Delta(K^{i})$, i.e.,
    $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j},$
    then it holds that there does not exist any $\beta\in\mathbb{R}$ such that:
    $\{V^{g^{j}}(\phi^{i}(k)\circ\pi)\}_{k=1}^{K^{i}}=\beta\mathbf{1}_{K^{i}}.$
@@
Let Assumptions C.2 and C.3 hold. Assume that $\pi_{n}\to\pi$ pointwise as $n\to\infty$, then the following is satisfied: [^4] Pointwise convergence of $\pi_{n}$ to $\pi$ ensures that for any $i\in\mathcal{N}$, $j\in[J]$, and $k\in[K^{i}]$: $V^{r^{i}}(\phi^{i}(k)\circ\pi_{n})\to V^{r^{i}}(\phi^{i}(k)\circ\pi)\quad\text{and}\quad V^{g^{j}}(\phi^{i}(k)\circ\pi_{n})\to V^{g^{j}}(\phi^{i}(k)\circ\pi)$ pointwise as $n\to\infty$.
@@
    Consider $rank(\psi)$ and $rank(\psi_{n})$: Since $\psi_{n}=\psi=\sum_{k=1}^{K^{i}}\alpha_{k}-1$, therefore:
@@
    Without loss of generality, we assume that $rank(\{\xi_{j}\}_{j\in[J]})=J$ since otherwise the redundant constraints can be removed. Note that removing redundant constraints from $LP^{i}(\pi)$ does not affect $LP^{i}_{n}$ since $I=\{m\in[J+K^{i}]\ \lvert\ \xi_{m}(\alpha)=0,\,\forall\alpha\in H(\xi,\psi)\}$ depends only on the limiting $LP^{i}(\pi)$. Next, we divide the satisfaction of each constraint $\xi_{j}$ for $j\in[J]$ into three distinct cases:
    1. (a)
      For all $\alpha\in\Delta(K^{i})$, we have $\xi_{j}(\alpha)=0$.
    2. (b)
      For all $\alpha\in\Delta(K^{i})$, we have $\xi_{j}(\alpha)\geq 0$. Furthermore, there exists a feasible $\alpha\in\Delta(K^{i})$ and a strictly infeasible solution $\alpha^{\prime}\in\Delta(K^{i})$, i.e., $\xi_{j}(\alpha)=0$ and $\xi_{j}(\alpha^{\prime})>0$, respectively.
    3. (c)
      There exists a strictly feasible solution $\alpha\in\Delta(K^{i})$ such that $\xi_{j}(\alpha)<0$, hence $j\notin I$.
    In case (a), the constraint can be removed from the linear program $LP^{i}(\pi)$ as it does not affect the feasible set. Without loss of generality, assume that there are $J_{1}\in[J]$ constraints $\{\xi_{j}\}$ for $j\in[J_{1}]$ satisfying condition (b), and the remaining $J-J_{1}$ constraints $\{\xi_{j}\}$ for $j=J_{1}+1,\dots,J$ satisfy condition (c). Therefore, we have $J_{1}]\subseteq I$.
    For the additional inequality constraints $\xi_{k}$, where $k=J+1,\dots,J+K^{i}$, Assumption C.3 guarantees the existence of a feasible $\alpha\in\Delta(K^{i})$ with positive weights, implying again that $I=[J_{1}]$.
    Suppose that $(\xi_{I},\psi)$ has full rank, i.e., $\mathrm{rank}(\xi_{I},\psi)=J_{1}+1$. Then, the following inequality holds:
    $\mathrm{rank}(\xi_{n,I},\psi_{n})\leq\mathrm{rank}(\xi_{I},\psi),\quad\forall n,$
    since $\mathrm{rank}(\xi_{n,I},\psi_{n})\leq J+1$. Hence, Inequality (28) is satisfied.
    Now, we show that $\mathrm{rank}(\xi_{I},\psi)=J_{1}+1$. Any linear mapping $(\xi_{I},\psi):\mathbb{R}^{K^{i}}\to\mathbb{R}^{J_{1}+1}$ can be represented as $\alpha^{\top}M-(1,c)^{\top}$, where $M\in\mathbb{R}^{(J_{1}+1)\times K^{i}}$ is defined as:
    $M:=\begin{bmatrix}1&1&\dots&1\\ V^{g^{1}}(\phi^{i}(1)\circ\pi)&V^{g^{1}}(\phi^{i}(2)\circ\pi)&\dots&V^{g^{1}}(\phi^{i}(K^{i})\circ\pi)\\ \vdots&\vdots&\ddots&\vdots\\ V^{g^{J_{1}}}(\phi^{i}(1)\circ\pi)&V^{g^{J_{1}}}(\phi^{i}(2)\circ\pi)&\dots&V^{g^{J_{1}}}(\phi^{i}(K^{i})\circ\pi)\end{bmatrix}$.
    By Assumption C.3, we have $\mathrm{rank}(M)=J_{1}+1$, which concludes the proof.
@@
    We need to verify that Assumption C.3 holds for all $\pi\in\mathcal{C}_{\pi}$ and all $i\in\mathcal{N}$.
    For any policy lying in the interior of $\mathcal{C}_{\pi}$, the identity modification $\phi^{i}(\mathrm{id})\in\Phi^{i}_{\mathrm{M,det}}$ always ensures $V^{g^{j}}(\phi^{i}(\mathrm{id})\circ\pi)>c^{j}$. Now, pick any $\alpha^{\prime}\in\Delta(K^{i})$ such that $\alpha^{\prime}_{k}>0$ for all $k\in[K^{i}]$, and define $\alpha^{\prime\prime}:=(1-\varepsilon)\mathbf{1}_{\mathrm{id}}+\varepsilon\alpha^{\prime}$, where $\mathbf{1}_{\mathrm{id}}$ is the canonical basis vector whose position of $1$ corresponds to the identity modification. By choosing $\varepsilon>0$ sufficiently small, it follows that $\sum_{k=1}^{K^{i}}\alpha^{\prime\prime}_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)>c^{j}$ with $\alpha^{\prime\prime}_{k}>0$ for all $k\in[K^{i}]$.
    For policies on the boundary of $\mathcal{C}_{\pi}$, if there exists a strictly infeasible $\alpha\in\Delta(K^{i})$ such that
    $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j}.$
    Then, there must exist at least one modification $\phi^{i}(k)$ for which $V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j}$. On the other hand, under the identity modification $\phi^{i}(\mathrm{id})$, we have $V^{g^{j}}(\phi^{i}(\mathrm{id})\circ\pi)=c^{j}$. Thus, there does not exist any $\beta\in\mathbb{R}$ such that $\{V^{g^{j}}(\phi^{i}(k)\circ\pi)\}_{k=1}^{K^{i}}=\beta\mathbf{1}_{K^{i}}$. Hence, together with Assumption 4.2, Assumption C.3 is satisfied.
@@
$\overset{(i)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}}\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\phi^{i}(k)\circ\pi}(s_{t},a_{t})g_{t}^{j}(s_{t},a_{t})$